OpenStudy (anonymous):

A 506g mass oscillates with an amplitude of 12.0cm on a spring whose spring constant is 23.0 N/m . Determine the period. Determine the maximum speed. Determine the total energy.

OpenStudy (anonymous):

T=2*pi*sqrt(m/k)=2 *pi*sqrt(0.506 / 23) = 0.93 s Total energy E=0.5*k*A=0.5*23*0.12=1.38 J Max speed is when kinetic energy T=E, so E=T=0.5mv^2 v=sqrt(2*E/m)=sqrt(2*1.38/0.506)= 2.33 m/s